OpenStudy (anonymous):

Is the function defined by x^2 - sin x + 5 continuous at x = pi?

OpenStudy (rational):

Hint : sum of two continuous functions is continuous

OpenStudy (rational):

you don't need to do any work if you're allowed to use that fact

OpenStudy (rational):

otherwise you may use LHL=RHL=f(pi) to prove continuity

OpenStudy (anonymous):

Sum of two continuous functions is continuous, how do I go down this method?

OpenStudy (rational):

x^2+5 is continuous as all polynomials are continuous -sinx is another familiar continuous function so the sum of these will be another continuous function

OpenStudy (rational):

based on your previous problems i don't think you're allowed to finish it off simply like that

OpenStudy (anonymous):

Polynomials and trig functions are both continuous? even exponents?

OpenStudy (rational):

All polynomials are continuous sinx and cosx are continuous you need to know them before using "sum of two continuous function... " thingy

OpenStudy (anonymous):

so f(x) = sin x + cos x f(x) = sin x - cos x f(x) = sinx . cos x Are all continuous just because they are trig functions?

OpenStudy (rational):

I never said all trig functions are continuous

OpenStudy (rational):

sinx and cosx are continuous so their sum, sinx+cosx is also continuous

OpenStudy (rational):

hey you don't need all this for exam on 16th

OpenStudy (rational):

i feel like your prof wants you to use LHL=RHL=f(pi) and finish off

OpenStudy (anonymous):

No its a question in my book, f(x) = sin x + cos x I have about 9 questions with trig functions in the piecewise type

OpenStudy (rational):

whats the complete question ?

OpenStudy (anonymous):

um f(x)= kx + 1, if x < or = pi cos x, if x > pi find k do you want me to post a new question for this?

OpenStudy (rational):

you may use LHL=RHL=f(a) for all these problems may be forget about "sum of two continuous functions if continuous" fact for now

OpenStudy (rational):

I presume you're given that f(x) is a continuous function ?

OpenStudy (rational):

``` f(x)= kx + 1, if x < or = pi cos x, if x > pi ```

OpenStudy (rational):

Since f(x) is continuous at x=pi, the left and right limits must be equal : \[ \large \lim\limits_{x\to \pi^{-}} f(x) =\lim\limits_{x\to \pi^{+}} f(x) \] \[\large k\pi+1 = \cos(\pi)\] solve \(k\)

OpenStudy (anonymous):

okay im a little stuck, can you solve this for me?

OpenStudy (rational):

Since f(x) is continuous at x=pi, the left and right limits must be equal : \[ \large \lim\limits_{x\to \pi^{-}} f(x) =\lim\limits_{x\to \pi^{+}} f(x) \] \[\large k\pi+1 = \cos(\pi)\] \[\large k\pi+1 = -1\] \[\large k\pi = -2\] \[\large k = -\dfrac{2}{\pi}\] done !!!

OpenStudy (rational):

Easy..

OpenStudy (anonymous):

That's why.. I put cos(pi) as 1..

OpenStudy (anonymous):

Got it, brb attempting all 8 questions under this mini-section

OpenStudy (rational):

looks all of them can be solved using LHL=RHL=f(a) trick just tag me if you feel stuck.. il be around mostly...